OpenStudy (anonymous):

Find all solutions (to three decimal places) of the equation in the given interval. ????????? cos^2 x - 2cos x - 1= 0 [0,π]

OpenStudy (anonymous):

remember double angle identity 2cos x - 1= cos(2A)

OpenStudy (anonymous):

cos(2A) = cos^2-sin^2

OpenStudy (anonymous):

cos^2 -cos^2 -sin^2 =0

OpenStudy (anonymous):

can you continue?

OpenStudy (anonymous):

uhh...im trying to find solutions, not solving a proof...

OpenStudy (phi):

how about let y = cos x then your equation is y^2 - 2y - 1 =0 find the solutions for y as a first step. what do you get ?

OpenStudy (phi):

use the quadratic formula http://www.purplemath.com/modules/solvquad4.htm

OpenStudy (anonymous):

I already got it.